For each pair of functions f and g below, find f(g(x)) and g(f(x)). Then, determine whether f and g are inverses of each other.
f(x)= 2/x, x is not equal to 0
g(x)= 2/x, x is not equal to 0
f(g(x))=
g(f(x))=
Are they inverses of each other? Thanks

Answer by stanbon(75887)

For each pair of functions f and g below, find f(g(x)) and g(f(x)). Then, determine whether f and g are inverses of each other.
f(x)= 2/x, x is not equal to 0
g(x)= 2/x, x is not equal to 0
f(g(x))= f{2/x) = 2/(2/x) = x
--------
g(f(x))= g(2/x) = 2/(2/x) = x
-----------------------------------
Are they inverses of each other?
Yes, because each of the functions undoes what
the other does.
Cheers,
Stan H.